4x-2=5x+7
4y+3=-2y-5
-7a+4=5a-2
-5x-7=-3x+2

Answer by checkley71(8403)

4X-2=5X+7 OR 4X-5X=7+2 OR -X=9 OR X=-9
4Y+3=-2Y-5 OR 4Y+2Y=-5-3 OR 6Y=-8 OR Y=-(8/6) OR Y=-(4/3) OR Y=-1 1/3
-7A+4=5A-2 OR -7A-5A=-2-4 OR -12A=-6 OR 12A=6 OR A=6/12 OR A=1/2
-5X-7=-3X+2 OR -5X+3X=+2+7 OR -2X=9 OR 2X=-9 OR X=-9/2 OR X=-4 1/2